# On Temperature Scaling and Conformal Prediction of Deep Classifiers

### v3

@@ Acknowledgment
We start with T $>$ 1 branch: for which $\hat{q}\geq\hat{q}_{T}$:
$\forall M\in[L]:\quad\sum_{i=1}^{M}\hat{\pi}_{i}-\sum_{i=1}^{M}\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\iff\forall M\in[L]:\quad\sum_{i=1}^{M}\hat{\pi}_{i}\geq\hat{q}-\hat{q}_{T}+\sum_{i=1}^{M}\hat{\pi}_{T,i}$
Note that for every $0<x\leq\hat{q}$ we have $M_{x}:=\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i}\geq x\}\leq L$. For $M_{x}>1$ (i.e., not minimal possible value), the event implies $\hat{q}-\hat{q}_{T}+\sum_{i=1}^{M_{x}-1}\hat{\pi}_{T,i}\leq\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}<x$. This implies that $\min\{l:\hat{q}-\hat{q}_{T}+\sum_{i=1}^{l}\hat{\pi}_{T,i}\geq x\}$ cannot be smaller than $M_{x}$. This also trivially holds for $M_{x}=1$ (the minimal set size). That is,
  $\displaystyle\Longrightarrow\forall\,\,0<x\leq\hat{q}:\quad\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i}\geq x\}\leq\min\{l:\hat{q}-\hat{q}_{T}+\sum_{i=1}^{l}\hat{\pi}_{T,i}\geq x\}$
Let us pick $x=\hat{q}$:
$\displaystyle\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i}\geq\hat{q}\}\leq\min\{l:\hat{q}-\hat{q}_{T}+\sum_{i=1}^{l}\hat{\pi}_{T,i}\geq\hat{q}\}$
$\displaystyle\iff\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i}\geq\hat{q}\}\leq\min\{l:\sum_{i=1}^{l}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}\iff L\leq L_{T}$
We continue with 0 $<$ T $<$ 1 branch:, for which $\hat{q}_{T}\geq\hat{q},$ we will take similar steps:
$\forall M\in[L]:\quad\sum_{i=1}^{M}\hat{\pi}_{i}-\sum_{i=1}^{M}\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\iff\forall M\in[L]:\quad\sum_{i=1}^{M}\hat{\pi}_{T,i}\geq\hat{q}_{T}-\hat{q}+\sum_{i=1}^{M}\hat{\pi}_{i}$
Note that for every $0<x\leq\hat{q}_{T}$ we have $M_{x}:=\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i,T}\geq x\}\leq L$. For $M_{x}>1$ (i.e., not minimal possible value), the event implies $\hat{q}_{T}-\hat{q}+\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}\leq\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i,T}<x$. This implies that $\min\{l:\hat{q}_{T}-\hat{q}+\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}\geq x\}$ cannot be smaller than $M_{x}$. That is,
  $\displaystyle\Longrightarrow\forall\,\,0<x\leq\hat{q}_{T}:\quad\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i,T}\geq x\}\leq\min\{l:\hat{q}_{T}-\hat{q}+\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}\geq x\}$
Let us pick $x=\hat{q}_{T}$:
$\displaystyle\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i,T}\geq\hat{q}_{T}\}\leq\min\{l:\hat{q}_{T}-\hat{q}+\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}\geq\hat{q}_{T}\}$
$\displaystyle\iff\min\{l:\sum_{i=1}^{l}\hat{\pi}_{i,T}\geq\hat{q}_{T}\}\leq\min\{l:\sum_{i=1}^{M_{x}-1}\hat{\pi}_{i}\geq\hat{q}\}\iff L_{T}\leq L$
@@
$&\text{If}\quad 0<T<1:\quad\Delta z>\max\left\{\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{T}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right\}\implies\nabla_{z_{1}}g(\mathbf{z};T,M)>0\\ &\text{If}\quad T>1:\quad\Delta z>\max\left\{\frac{T}{T-1}\ln(4T),\frac{T-1}{T+1}\ln(4T(C-1)^{2})\right\}\implies\nabla_{z_{1}}g(\mathbf{z};T,M)<0\end{aligned}\right.$ (10)
@@
$\displaystyle\Delta z>\max\left(\frac{T}{T-1}\ln(4T),\frac{T-1}{T+1}\ln(4T(C-1)^{2})\right)\Longrightarrow\nabla_{z_{1}}g_{z}(\mathbf{z};T,M)<0$
@@
$\displaystyle\Delta z>\max\left(\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{T-1}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right)\Longrightarrow\nabla_{z_{1}}g_{z}(\mathbf{z};T,M)>0$
@@ A.1 On the link between (3) and (5)
With our assumption that $\hat{q}$ and $\hat{q}_{T}$ are associated with the same quantile sample $\mathbf{z}^{q}$, we have that $\hat{q}-\hat{q}_{T}$ in (3) can be written as $g(\mathbf{z}^{q};T,L^{q})$, where $L^{q}$ denotes the rank of the true label of $\mathbf{z}^{q}$. The left-hand side in (3) can be written as $g(\mathbf{z};T,M)$. Thus, it would facilitate the analysis to study the relation between $g(\mathbf{z};T,M)$ and $g(\mathbf{z}^{q};T,M)$, as in (5), rather than by $g(\mathbf{z};T,M)$ and $g(\mathbf{z}^{q};T,L^{q})$. Since $\Delta z^{q}=z^{q}_{(1)}-z^{q}_{(2)}\gg 1$, we have found both empirically and theoretically that $g(\mathbf{z}^{q},T,M)\approx g(\mathbf{z}^{q},T,L^{q})$, which justifies studying (5).
A.1.1 Empirical justification
Here, we demonstrate empirically that the difference $\left|g(\mathbf{z}^{q};T,M)-g(\mathbf{z}^{q};T,L^{q})\right|$ is negligible compared to the difference $\left|g(\mathbf{z};T,M)-g(\mathbf{z}^{q};T,M)\right|$.
We consider the quantile sample $\mathbf{z}^{q}$ associated with $0.9\displaystyle 1-\alpha=$ 0.9 and $L^{q}$ as the rank of its true label. Similarly to the analysis in Section 4, we treat a “typical” sample $\mathbf{z}$ as one with the median score, for which $L$ is chosen by the number of sorted softmax entries require for exceeding the score of $\mathbf{z}^{q}$.
For the CIFAR100-ResNet50 pair we get:
$\displaystyle\max_{M\in[L],T\in[0.5,5]}\left|g(\mathbf{z}^{q};T,M)-g(\mathbf{z}^{q};T,L^{q})\right|=0.0031,$
$\displaystyle\min_{M\in[L],T\in[0.5,5]}\left|g(\mathbf{z};T,M)-g(\mathbf{z}^{q};T,M)\right|=0.0513.$
For the ImageNet-ViT pair we get:
$\displaystyle\max_{M\in[L],T\in[0.5,5]}\left|g(\mathbf{z}^{q};T,M)-g(\mathbf{z}^{q};T,L^{q})\right|=0.051,$
$\displaystyle\min_{M\in[L],T\in[0.5,5]}\left|g(\mathbf{z};T,M)-g(\mathbf{z}^{q};T,M)\right|=0.2597.$
The difference $\left|g(\mathbf{z}^{q};T,M)-g(\mathbf{z}^{q};T,L^{q})\right|$ is more than an order of magnitude smaller than $\left|g(\mathbf{z};T,M)-g(\mathbf{z}^{q};T,M)\right|$. Similar results are obtained for the other pairs as well. This observation reinforces the study of $\left|g(\mathbf{z};T,M)-g(\mathbf{z}^{q};T,M)\right|$.
A.1.2 Theoretical Justification
Below, in Proposition A.6 we show that $\forall M,L^{q}\in[C]$ the difference $\left|g(\mathbf{z}^{q};T,M)-g(\mathbf{z}^{q};T,L^{q})\right|$ decays exponentially with $\Delta z^{q}$. As demonstrated in Figure 2, $\Delta z^{q}$ of the quantile sample is very high, which indicates the small value of the bound. The proof of the proposition appears in Section A.2.
@@ B.3.1 ECE vs NLL minimization
B.4 Metrics definitions
In section 3.1, we use metrics to represent average prediction set size, marginal coverage and conditional coverage. Here, we present the formulas for these metrics. Note that similar metrics have been used in (Ding et al., 2023; Angelopoulos et al., 2021).
We report metrics over the validation set which we denote by $\{(x_{i}^{(val)},y_{i}^{(val)})\}_{i=1}^{N^{(val)}}$, comprising of the samples that were not included in the calibration set or CP set. The metrics are as follows.
  • Average set size (AvgSize) – The mean prediction set size of the CP algorithm:
    $\text{AvgSize}=\frac{1}{N^{(val)}}\sum_{i=1}^{N^{(val)}}|C(x_{i}^{(val)})|.$ (16)
  • Marginal coverage gap (MarCovGAP) – The deviation of the marginal coverage from the desired $1-\alpha$:
    $\text{MarCovGap}=\left|\frac{1}{N^{(val)}}\sum_{i=1}^{N^{(val)}}\mathbbm{1}\{y_{i}^{(val)}\in C(x_{i}^{(val)})\}-(1-\alpha)\right|.$ (17)
  • Top-5% class-coverage gap (TopCovGap) – The deviation from the desired $1-\alpha$ coverage, averaged over the 5% of classes with the highest deviation:
[…]
B.4.1 Reliability diagrams
@@
ImageNet, ResNet152   CIFAR-100, ResNet50   CIFAR-10, ResNet50
@@
B.5 The effect of TS calibration on CP methods
@@ B.6.2 TopCovGap metric
B.6.3 CP threshold value $\hat{q}$
@@
B.6.4 The impact of TS at extremely low temperatures
@@ Appendix C Approximating the Trade-off via the Calibration Set
Unlike the curves in Figure 1, the curves in Figure 17 are not averaged over 100 trials but are based on a single trial. Due to randomization, the curves will vary between runs, so to better reflect the practitioner’s experience, we present results from 3 separate runs. In each of the runs in Figure 17 the marginal coverage is preserved and the curves of AvgSize and TopCovGap closely resemble the averaged ones shown in Figure 1, demonstrating the user’s ability to select $\hat{T}$ based on these single-trial graphs generated using small amount of data. Additionally, note that the procedure required to produce these approximated curves is executed offline during calibration and has a negligible runtime compared to the offline training of DNNs.

### v4

@@ Acknowledgment
We start with T $>$ 1 branch, for which $\hat{q}\geq\hat{q}_{T}$:
$\displaystyle\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}\geq\hat{q}-\hat{q}_{T}\iff\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{i}\geq\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}+\hat{q}-\hat{q}_{T}$ (10)
Note that $L:=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{i}\geq\hat{q}\}\implies\sum_{i=1}^{L-1}\hat{\pi}_{i}<\hat{q}$. Thus, if (10) holds, then:
$\displaystyle\hat{q}>\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}+\hat{q}-\hat{q}_{T}\iff\hat{q}_{T}>\sum\nolimits_{i=1}^{L-1}\hat{\pi}_{T,i}.$
Therefore, $L_{T}:=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}$ cannot be smaller than $L$.
We continue with 0 $<$ T $<$ 1 branch, for which $\hat{q}_{T}\geq\hat{q},$ we will take similar steps:
$\displaystyle\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}-\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{T,i}\leq\hat{q}-\hat{q}_{T}\iff\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{T,i}\geq\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}+\hat{q}_{T}-\hat{q}$ (11)
Note that $L_{T}:=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{T,i}\geq\hat{q}_{T}\}\implies\sum_{i=1}^{L_{T}-1}\hat{\pi}_{T,i}<\hat{q}_{T}$. Thus, if (11) holds, then:
$\displaystyle\hat{q}_{T}>\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}+\hat{q}_{T}-\hat{q}\iff\hat{q}>\sum\nolimits_{i=1}^{L_{T}-1}\hat{\pi}_{i}.$
Therefore, $L:=\min\{\ell:\sum\nolimits_{i=1}^{\ell}\hat{\pi}_{i}\geq\hat{q}\}$ cannot be smaller than $L_{T}$.
@@
$&\text{If}\quad 0<T<1:\quad\Delta z>\max\left\{\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{T}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right\}\implies\nabla_{z_{1}}g(\mathbf{z};T,M)>0\\ &\text{If}\quad T>1:\quad\Delta z>\max\left\{\frac{T}{T-1}\ln(4T),\frac{{T}}{T+1}\ln(4T(C-1)^{2})\right\}\implies\nabla_{z_{1}}g(\mathbf{z};T,M)<0\end{aligned}\right.$ (12)
@@
$\displaystyle\Delta z>\max\left(\frac{T}{T-1}\ln(4T),\frac{{T}}{T+1}\ln(4T(C-1)^{2})\right)\Longrightarrow\nabla_{z_{1}}g_{z}(\mathbf{z};T,M)<0$
@@
$\displaystyle\Delta z>\max\left(\frac{T}{T-1}\ln\left(\frac{T}{4}\right),\frac{{T}}{T+1}\ln\left(\frac{4(C-1)^{2}}{T}\right)\right)\Longrightarrow\nabla_{z_{1}}g_{z}(\mathbf{z};T,M)>0$
@@ A.1 On the link between (3) and (5)
With our assumption that $\hat{q}$ and $\hat{q}_{T}$ are associated with the same quantile sample $\mathbf{z}^{q}$, we have that $\hat{q}-\hat{q}_{T}$ in (3) can be written as $g(\mathbf{z}^{q};T,L^{q})$, where $L^{q}$ denotes the rank of the true label of $\mathbf{z}^{q}$. The left-hand side in (3) can be written as $g(\mathbf{z};T,{L_{T}-1})$ for $T<1$ and $g(\mathbf{z};T,{L-1})$ for $T>1$. In this section, we theoretically show the similarity between $g(\mathbf{z};T,M)$ and $g(\mathbf{z};T,L_{q})$ $\forall M\in[C]$. Specifically, we prove exponential decay of this distance as function of $\Delta z$. According to Figure 2, this justifies the assumption of $g(\mathbf{z}^{q},T,M)\approx g(\mathbf{z}^{q},T,L^{q})$, which leads to studying (5).
@@ B.3.1 ECE vs NLL minimization
B.3.2 Reliability diagrams
@@
ImageNet, DenseNet121   CIFAR-100, DenseNet121   CIFAR-10, ResNet34
@@
B.4 The effect of TS calibration on CP methods
@@ B.5.2 TopCovGap metric
B.5.3 AvgCovGap metric
In this subsection, we define and examine the average class-coverage gap (AvgCovGap) – a metric that measures the deviation from the target $1-\alpha$ coverage, averaged across all classes:
$\text{AvgCovGap}=\frac{1}{C}\sum_{y=1}^{C}\left|\frac{1}{|I_{y}|}\sum_{i\in I_{y}}\mathbbm{1}\left\{y_{i}^{(val)}\in C\left(x_{i}^{(val)}\right)\right\}-(1-\alpha)\right|,$
where $I_{y}=\{i\in[N_{val}]:y_{i}^{(val)}=y$}. The following experimental results for AvgCovGap closely mirror those observed with the TopCovGap metric.
ImageNet, ViT-B/16   CIFAR-100, ResNet50   CIFAR-10, ResNet50
ImageNet, ResNet152   CIFAR-100, DenseNet121   CIFAR-10, ResNet34
Figure 14: Conditional Coverage Metric. AvgCovGap using LAC, APS and RAPS with $\alpha=0.1$ and CP set size $10\%$ versus the temperature $T$ for additional dataset-model pairs.
ImageNet, ViT-B/16   CIFAR-100, ResNet50   CIFAR-10, ResNet50
ImageNet, ResNet152   CIFAR-100, DenseNet121   CIFAR-10, ResNet34
Figure 15: Conditional Coverage Metric. AvgCovGap using LAC, APS and RAPS with $\alpha=0.05$ and CP set size $10\%$ versus the temperature $T$ for additional dataset-model pairs.
B.5.4 CP threshold value $\hat{q}$
@@
B.5.5 The impact of TS at extremely low temperatures
@@
B.6 Illustrating the similarity between $\mathbf{z}^{q}$ and $\mathbf{z}^{q}_{T}$
Figure 2 in Section 4.2 shows a strong correlation between the APS score and $\Delta z:=z_{(1)}-z_{(2)}$, indicating that quantile samples typically have a dominant entry in their logit vectors ($z_{(1)}^{q}\gg z_{(2)}^{q}$). This behavior persists under TS, i.e., for the logits vector $\mathbf{z}^{q}_{T}$ of the quantile sample when TS has been applied with some temperature $T$. Recall that the softmax operation applied to the logits vectors is invariant to constant shifts and further amplifies the dominance of larger entries ($\pi_{i}\propto\exp(z_{i})$). Thus, having $z^{q}_{T,(1)}-z^{q}_{T,(2)}\approx z^{q}_{(1)}-z^{q}_{(2)}\gg 1$, practically implies that $\mathbf{z}^{q}_{T}\approx\mathbf{z}^{q}$.
Here, we demonstrate this similarity using concrete examples of the largest 5 elements of these vector for various dataset–model pairs and temperature settings. To simplify the comparison, we report the values after applying softmax. That is, we report the values of $\bm{\sigma}(\mathbf{z}^{q}_{T})$.
CIFAR10-ResNet34:
$\bm{\sigma}(\mathbf{z}^{q})[:5]=[9.9997\mathrm{e-}01,8.2661\mathrm{e-}06,7.3281\mathrm{e-}06,6.7146\mathrm{e-}06,2.0070\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=T^{*}})[:5]=[9.9997\mathrm{e-}01,1.4801\mathrm{e-}05,8.0310\mathrm{e-}06,2.0299\mathrm{e-}06,1.6542\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=0.5})[:5]=[9.9998\mathrm{e-}01,1.1870\mathrm{e-}05,4.1955\mathrm{e-}06,2.0069\mathrm{e-}06,1.9213\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=2})[:5]=[9.9997\mathrm{e-}01,2.0271\mathrm{e-}05,2.0588\mathrm{e-}06,1.7960\mathrm{e-}06,1.6457\mathrm{e-}06]$
CIFAR100-DenseNet121:
$\bm{\sigma}(\mathbf{z}^{q})[:5]=[9.9993\mathrm{e-}01,5.6212\mathrm{e-}05,5.1748\mathrm{e-}06,3.4063\mathrm{e-}06,1.2153\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=T^{*}})[:5]=[9.9992\mathrm{e-}01,4.9569\mathrm{e-}05,8.1525\mathrm{e-}06,4.1283\mathrm{e-}06,1.8294\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=0.5})[:5]=[9.9997\mathrm{e-}01,1.2354\mathrm{e-}06,6.5692\mathrm{e-}07,4.6121\mathrm{e-}07,4.4120\mathrm{e-}07]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=2})[:5]=[9.9996\mathrm{e-}01,7.2431\mathrm{e-}06,6.0478\mathrm{e-}06,4.3499\mathrm{e-}06,3.7969\mathrm{e-}06]$
ImageNet-ResNet152:
$\bm{\sigma}(\mathbf{z}^{q})[:5]=[9.9991\mathrm{e-}01,5.8738\mathrm{e-}05,1.5782\mathrm{e-}05,8.8207\mathrm{e-}06,4.0592\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=T^{*}})[:5]=[9.9992\mathrm{e-}01,3.1237\mathrm{e-}05,9.8604\mathrm{e-}06,9.4194\mathrm{e-}06,3.6732\mathrm{e-}06]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=0.5})[:5]=[9.9993\mathrm{e-}01,3.4423\mathrm{e-}05,2.5229\mathrm{e-}05,1.9685\mathrm{e-}05,1.9673\mathrm{e-}05]$
$\bm{\sigma}(\mathbf{z}^{q}_{T=2})[:5]=[9.9989\mathrm{e-}01,2.9378\mathrm{e-}04,6.1594\mathrm{e-}06,1.5067\mathrm{e-}06,6.1018\mathrm{e-}07]$
Across all these dataset–model pairs and temperature settings, the quantile samples exhibit strong similarity. This observation further supports the technical assumption in in Section 4.2 that both $\hat{q}$ and $\hat{q}_{T}$ correspond to the same underlying sample.
@@ Appendix C Approximating the Trade-off via the Calibration Set
Unlike the curves in Figure 1, the curves in Figure 19 are not averaged over 100 trials but are based on a single trial. Due to randomization, the curves will vary between runs, so to better reflect the practitioner’s experience, we present results from 3 separate runs. In each of the runs in Figure 19 the marginal coverage is preserved and the curves of AvgSize and TopCovGap closely resemble the averaged ones shown in Figure 1, demonstrating the user’s ability to select $\hat{T}$ based on these single-trial graphs generated using small amount of data. Additionally, note that the procedure required to produce these approximated curves is executed offline during calibration and has a negligible runtime.